# Chapter 1      Introduction to IoT

### 09 Dec 2023 v1

@@ 1.6.2 Medical and Healthcare
ECG Monitoring
An Electrocardiogram (ECG) measures the heart’s electrical signals and serves as an indicator of heart health, aiding in the detection of conditions like arrhythmia, prolonged QT interval, and myocardial ischemia. IoT integration in ECG monitoring has enabled early detection of heart abnormalities.
Wu et al. [82] have introduced an ECG data monitoring system utilizing a bipotential chip attached to the user’s t-shirt, transmitting data to a smartphone via Bluetooth. Real-time ECG data monitoring becomes feasible when IoT systems are combined with big data analytics. Bansals and Gandhi [83] have presented a scheme that amalgamated nanoelectronics, IoT, and big data analytics for long-term and continuous ECG data monitoring.
@@
Diabetes is a medical condition characterized by higher blood glucose levels than those found in individuals without diabetes. Various types of diabetes exist, and doctors use three primary methods to identify them: random plasma glucose tests, fasting plasma glucose tests, and oral glucose tolerance tests. Nevertheless, the fingerstick method, involving a small pinprick to the fingertip followed by blood glucose level measurement, remains the most commonly used diagnostic approach.
The advent of IoT technology has brought about improvements in this process, making it quicker and more convenient for patients. Istepanian et al. [84] have introduced an IoT-integrated noninvasive blood glucose monitoring device to continuously monitor glucose levels, eliminating the need for fingerstick testing. It is worth noting that optical sensors, such as infrared LED and near-infrared photodiode setups, have also been employed for glucose level measurements. In these approaches, the reflected light signal from the human body is used to calculate the glucose level [85].
@@
Monitoring body temperature is crucial for maintaining homeostasis and plays a pivotal role in various diagnostic processes. Traditionally, temperature measurement methods involve using thermometers placed in the mouth, ear, or rectum, but these methods often cause discomfort for the patient and pose an elevated risk of infection. However, recent advancements in IoT-based temperature monitoring applications have effectively addressed these issues.
For instance, one innovative approach by [86] where authors have introduced a 3D printed wearable device designed to be inserted into the ear. This device utilizes an infrared sensor to measure temperature from the tympanic membrane, ensuring accuracy while remaining unaffected by external factors like environmental conditions or physical activities. In another significant development, Gunawan [87] has created an IoT-based temperature monitoring system employing both Arduino and Raspberry Pi. This system records temperature data into a database, which users can conveniently access through a web page via their mobile phones or computers.
@@
In many diagnostic processes, measuring blood pressure is a compulsory step. However, the major issue with the traditional method is that it requires one person to record the blood pressure. Therefore, the integration of IoT into blood pressure monitoring has been a blessing for both doctors and patients. In [88], a wearable cuffless gadget has been proposed, capable of measuring both systolic and diastolic pressure, with the results stored in the cloud. A cloud and fog-based architecture have been proposed in [89] for long-term, real-time blood pressure monitoring, with the data also being stored for future references.
@@
Pulse oximetry, a highly beneficial noninvasive device for measuring oxygen saturation, addresses the limitations of traditional methods and allows for real-time monitoring. The integration of IoT-based technology has led to significant advancements in pulse oximetry, particularly in the healthcare industry. In the he study [90], an advanced noninvasive pulse oximetry system has been proposed, capable of measuring oxygen levels, heart rate, and pulse parameters, while transmitting this data to a central server. Another study [91], introduced an integrated alert system that sends notifications when oxygen levels reach a critical threshold. This system incorporates a pulse oximeter and WLAN router connected through the Blynk server.
@@
Monitoring a person’s mood is crucial for understanding their mental state and promoting a mentally healthy life. It also aids healthcare specialists and psychologists in treating various mental health conditions such as depression, stress, and bipolar disorder. Self-monitoring emotional stability empowers individuals to proactively address their mental well-being.
The integration of IoT into this domain offers numerous advantages. Continuous remote monitoring becomes possible, greatly benefiting individuals with severe mental illnesses by enabling real-time responses to emergencies. Furthermore, IoT can detect a person’s mental state by analyzing heartbeats through wearable devices. Kaur et al. [92] have proposed a wearable device capable of tracking a driver’s emotions, including anger, stress, terror, and sadness. The intelligent system, by analyzing emotion variations, determines whether the driver has entered a subconscious state and stops the vehicle’s DC motor accordingly. In another study, machine learning and IoT were combined to monitor a user’s mood by analyzing their social networking activities and heart rate through a wearable device [93].
@@
Adhering to a prescribed medication schedule is crucial for maintaining one’s health. However, this can be particularly challenging for elderly individuals who may experience memory issues, dementia, or cognitive decline. Managing daily medication routines becomes increasingly difficult for them. Fortunately, the integration of IoT offers a solution to this problem, and numerous research efforts have explored using IoT to track patients’ medication compliance.
In one study [94], a medical box was developed with the primary function of reminding individuals to take their medications. Notably, this system features three separate trays for different times of the day (morning, afternoon, and night). Additionally, it can measure vital health parameters such as blood pressure, temperature, blood oxygen levels, and ECG readings. A mobile application was designed to facilitate two-way communication between patients and doctors, and the recorded information is accessible to both parties. Furthermore, authors in [95] have proposed storage conditions, including temperature and humidity solutions, to help users maintain an optimal environment for their medications.
@@
Wheelchairs are vital tools in the lives of individuals who are physically unable to move independently, providing both physiological assistance and psychological support. However, for individuals with brain damage who lack the capability to operate a wheelchair, researchers have been exploring the addition of navigation and tracking systems, with IoT playing a crucial role in these wheelchair advancements.
In one study [96], an IoT-based steering and real-time obstacle detection system for wheelchairs has been proposed. This system can identify obstacles through image processing methods, enhancing the wheelchair’s safety and usability. An advanced and automated smart wheelchair was reported in another study [97]. This wheelchair not only monitored movement but also offered features such as an umbrella, foot mat, head mat, and obstacle detection. These innovations significantly improved interaction with the living environment and enhanced the user’s overall experience.
@@
Rehabilitation is a critical aspect of restoring an individual’s functional abilities, aiming to help those with diseases or disabilities regain their normal lifestyle. The application of IoT in this field is versatile and has proven effective in various areas, including cancer treatment, sports injury recovery, stroke rehabilitation, and addressing physical disabilities.
For instance, an innovative smart walker was introduced in a particular study [98]. This walker utilizes a multi-modal sensor to monitor the patient’s walking pattern, and the collected data can be accessed by doctors and caregivers through a mobile application, facilitating better monitoring and support. A sports rehabilitation system was reported which monitors various parameters such as temperature, motion posture, electromyography (EMG), electrocardiography (ECG), among others, and provides real-time feedback to athletes [99]. These examples highlight the potential of IoT integration in modernizing rehabilitation systems, making them more efficient and effective.
@@
The quality of an individual’s life is significantly influenced by their level of fitness and daily physical activity. To facilitate fitness monitoring and promote healthier lifestyles, various applications leveraging the IoT have been developed. One approach involves utilizing smartphone accelerometer data to assess users’ activity levels. This data can reveal metrics such as the number of steps taken, the duration of physical activity, and periods of inactivity. Wearable fitness trackers are readily available in the market and have gained popularity as convenient devices for monitoring fitness levels. Additionally, IoT sensors can be integrated with gym equipment to provide real-time monitoring of individuals’ fitness progress. For instance, in a notable study [100], a “smart mat” have been introduced. This mat is designed to measure the number of exercise steps performed on its surface, offering valuable insights into users’ workout routines.
In the field of sports science, fitness assessment and training load monitoring have proven to be invaluable tools for both coaches and athletes. IoT technology enables the continuous and non-invasive monitoring of physiological biomarkers, such as saliva or sweat. This data can be leveraged to optimize hydration strategies, ultimately enhancing athletes’ overall performance [101]. These IoT-driven fitness monitoring applications and devices contribute to a more proactive approach to health and fitness, promoting active lifestyles and well-being.
@@
The application of IoT in the healthcare industry is incredibly diverse, extending far beyond the previously mentioned areas. There are numerous domains where IoT has already been implemented and where its potential benefits are being realized, leading to a significant increase in the adoption of Healthcare IoT (HIoT) technology.
For instance, in cancer treatment, IoT-based methods have emerged as powerful tools. An innovative IoT-based cancer treatment approach is introduced in a recent study, that encompasses various stages, including chemotherapy and radiotherapy [102]. Additionally, this system securely stores lab test results on a cloud server, allowing physicians to monitor medication dosages and enabling remote consultations through a dedicated mobile application. Furthermore, HIoT has found applications in detecting skin lesions [103], with notable advancements in lung cancer detection through state-of-the-art machine learning algorithms and IoT-based systems [104, 105, 106].
In the realm of surgical training and medical procedures, IoT has been harnessed to create next-generation solutions. One such development involves a surgical training framework that employs virtual reality to simulate realistic training environments. This framework also enables interaction with surgeons from around the world, fostering collaborative learning and expertise sharing [107]. Additionally, a human-robot collaborative system has been devised for performing minimally invasive surgeries [108]. Monitoring hemoglobin levels in the blood has become more accessible through portable devices equipped with photoplethysmography (PPG) sensors, light-emitting diodes (LEDs), and photodiodes. These devices enable the non-invasive measurement of hemoglobin levels, enhancing healthcare monitoring and diagnosis [109]. Asthma Monitoring is another notable sector where sensor-based asthma monitoring system is used to track the user’s respiratory rate [110].
Numerous other HIoT applications are currently in use or under research, underscoring the ongoing revolutionary impact of IoT in the field of healthcare. It is expected that IoT will continue to drive advancements and improvements in healthcare delivery and patient outcomes.
@@ 1.6.3 Smart Agriculture and Environment
Agriculture holds a crucial position in a country’s economic progress. However, crop yields can be significantly impacted by various factors, including soil moisture and environmental variables like carbon dioxide levels, temperature, and humidity. To enhance agricultural outcomes, it becomes essential to implement robust surveillance systems in the fields. Achieving this efficiently is made possible through the integration of the IoT. The following section has explored several applications in this agricultural domain as well as a summary of the related research work in Table 1.6.
@@
Applications Related Works
Water-saving irrigation [111, 112]
Crop growth environment monitoring [113, 114]
Animal and plant life information monitoring [115, 116]
Intelligent agricultural machinery [117, 118]
Agricultural product quality safety and traceability [119, 120]
Diseases Monitoring [121, 122]
@@
Water is a crucial factor in crop growth, yet its scarcity is an escalating concern. Different crops have varying water requirements, necessitating a dynamic irrigation approach. The integration of IoT technology in agriculture addresses this challenge by revolutionizing traditional flood irrigation methods, ultimately resolving water shortage problems.
To address these issues, Yang et al. [111] have proposed an intelligent system leveraging neural networks for water-efficient irrigation. This method enhances irrigation efficiency by minimizing human intervention and reducing wastage due to excessive drainage. Li et al. [112] have developed a series of GPRS DTU-based intelligent agricultural irrigation control systems capable of tailoring irrigation to crop growth needs.
@@
Various environmental factors, including temperature, humidity, air pressure, carbon dioxide levels, soil temperature, and soil pH, play a crucial role in crop growth. IoT devices integrated into agricultural systems can sense and analyze these environmental factors, enabling remote field monitoring and the creation of an optimal farming environment tailored to these variables.
Chen et al. [113] have proposed an environmental monitoring solution that utilizes WSNs for millet farmland. In contrast, Lin et al. [114] have designed a wireless environmental monitoring system that harnesses soil energy to enable cost-effective remote monitoring of farmland environments.
@@
Efficient agricultural production necessitates the monitoring of both animal and plant information, a pivotal phase in the process. This monitoring plays a substantial role in advancing improved production, increasing profitability, and ensuring high-quality product development.
@@
    Animal Life Information Monitoring
    Monitoring diverse aspects of animal behavior, including their food consumption, body temperature, activity levels, and health status, enables the tracking of their physiological and nutritional well-being, ensuring their healthy development. Similar to other monitoring systems, animal information monitoring includes various types of sensors, actuators, wireless connections, and terminal equipment.
    In [115], the authors have proposed an infrared-based body temperature measurement system for pigs, allowing the early detection of diseases .
@@
    Plant Life Information Monitoring
    Plant life information can be assessed through two distinct methods: external information and internal information. External data involves analyzing visual factors such as diseases, pests, and leaf color, while internal information includes parameters like chlorophyll content and photosynthetic rate. Various wireless sensor devices can be connected to plants to facilitate remote and continuous monitoring of their information, ultimately allowing for the early detection of diseases and promoting healthy plant growth.
    Porto et al. [116] have introduced a system for citrus traceability that examines the environmental conditions suitable for the optimal growth of citrus. This system is capable of identifying and preventing plant diseases, ensuring the robust health of citrus crops.
@@
Intelligent machinery demonstrates the capacity to autonomously manage a wide range of agricultural operations, encompassing tasks such as cultivation, sowing, transplanting, fertilization, pesticide application, feeding, irrigation, picking, and harvesting, all executed with precision and efficiency. Moreover, it has the capability to gather a variety of data about the farm, including soil moisture and water quality, as well as ambient information like temperature and humidity. These data resources can be effectively harnessed to implement precision agriculture and enhance breeding practices [123]. IoT technology plays a pivotal role in reducing manual labor in agricultural machinery operations. It enables remote monitoring and standardizes machinery functions through various technologies such as sensors and wireless communication methods.
Sowjanya et al. have introduced a versatile autonomous robot vehicle in [117] where the vehicle is equipped with Bluetooth technology for remote control and is capable of independently executing a range of tasks including farming, seeding, and irrigation. In a similar vein, researchers in [118] have presented a robot system designed for fruit detection. This system employs a single-lens multi-box detection method and a stereo camera to identify the position of fruits in three dimensions, enabling efficient fruit picking.
@@
IoT plays a pivotal role in enhancing agricultural product quality, safety, and traceability, particularly in the domains of product warehousing, logistics, and distribution. It enables automatic product identification, tracking, and the accurate tallying of incoming and outgoing items within warehouses. By leveraging WSNs, real-time remote monitoring of vehicles or containers used for product delivery can be achieved, ensuring the traceability of agricultural product origins and destinations. Numerous countries have recognized the imperative for real-time traceability and have developed intelligent systems to address this need. These systems include the American agricultural product traceability system, the European beef traceability system, the Swedish agricultural traceability system, the Japanese food traceability system, and the Australian livestock traceability system.
In the pursuit of agricultural product safety and traceability, Jiang et al. [119] have established a comprehensive agricultural product safety traceability platform. This platform enables the real-time automatic collection, processing, and display of agricultural product data, thereby enhancing product traceability and reducing the costs associated with tracking and monitoring agricultural items. Gu et al. have introduced a system in [120] designed to trace the complete lifecycle of pigs, covering stages such as slaughter, processing, and sales, thereby ensuring transparency and traceability throughout the pork production process.
@@
The integration of IoT for real-time and continuous monitoring offers farmers and relevant authorities the capability to identify diseases at an early stage and implement preventive measures before they escalate. IoT, when combined with machine learning, artificial intelligence, or image processing, has yielded various systems for early disease detection and prevention in agriculture.
The farm’s environment plays a pivotal role in disease occurrence, and in [121], authors have presented a framework that combines various sensor devices through wireless sensor networks for monitoring various environmental factors. This system collects real-time data and employs multiple machine learning models like KNN, Random Forest, Logistic Regression, and Linear Regression to analyze insights from this data. Additionally, a framework for environmental observation-based crop disease monitoring, which continuously monitors plant growth and diseases and notifies farmers when necessary, has been proposed in [122].

### 12 Jan 2024 v2

@@ 1.6.1 Smart Cities
1.6.2 Medical and Healthcare
The healthcare sector has experienced remarkable advancements through the integration of IoT, offering solutions to real-life healthcare challenges and enhancing people’s lifestyles. Researchers have proposed various applications in healthcare that utilize wearable sensor devices to monitor patients’ health, diagnose diseases, issue emergency alerts, and notify users when necessary. Remote monitoring saves time for both patients and doctors while reducing overall healthcare costs. Furthermore, sharing data collected from these wearable devices with healthcare researchers contributes to the development of safer and more timely healthcare solutions and aids in the discovery of cures and vaccines for emerging diseases. This section explores several IoT applications in healthcare, and a summary of the associated research papers is presented in Table 1.6.
Electrocardiogram (ECG) Monitoring
An ECG measures the heart’s electrical signals and serves as an indicator of heart health, aiding in the detection of conditions like arrhythmia, prolonged QT interval, and myocardial ischemia. An interesting example would be Wu et al.’s ECG data monitoring system [82], where a bipotential chip is utilized by attaching to the user’s t-shirt and transmitting data to a smartphone via Bluetooth. Combining IoT systems with big data analytics enables real-time ECG data monitoring.
@@ 1.6.2 Medical and Healthcare
Diabetes is a medical condition characterized by higher blood glucose levels than those found in individuals without diabetes. Among various approaches to identifying diabetes, the fingerstick method, involving a small pinprick to the fingertip followed by blood glucose level measurement, remains the most commonly used diagnostic approach. The advent of IoT technology has brought about improvements in this process, making it quicker and more convenient for patients. Istepanian et al. [84] have introduced an IoT-integrated noninvasive blood glucose monitoring device to continuously monitor glucose levels, eliminating the need for fingerstick testing. It is worth noting that optical sensors, such as infrared LED and near-infrared photodiode setups, have also been used for glucose level measurements.
@@
Traditionally, temperature measurement methods involve using thermometers placed in the mouth, ear, or rectum, but these methods often cause discomfort for the patient and pose an elevated risk of infection. However, recent advancements in IoT-based temperature monitoring applications have effectively addressed these issues. For instance, authors in [86] have introduced a 3D-printed wearable device designed to be inserted into the ear. This device utilizes an infrared sensor to measure temperature from the tympanic membrane, ensuring accuracy while remaining environment-independent.
@@
In many diagnostic processes, measuring blood pressure is a compulsory step. However, the major issue with the traditional method is that it requires one person to record the blood pressure. Therefore, the integration of IoT into blood pressure monitoring has been a blessing for both doctors and patients. A wearable cuffless gadget [88], for example, is capable of measuring both systolic and diastolic pressure, with the results stored in the cloud.
@@
Pulse oximetry, a highly beneficial noninvasive device for measuring oxygen saturation, addresses the limitations of traditional methods and allows for real-time monitoring. The integration of IoT-based technology has led to significant advancements in pulse oximetry, particularly in the healthcare industry. In a study [90], an advanced noninvasive pulse oximetry system has been proposed, capable of measuring oxygen levels, heart rate, and pulse parameters while transmitting this data to a central server.
@@
The integration of IoT into the mood monitoring domain offers numerous advantages. IoT can detect a person’s mental state by analyzing heartbeats through wearable devices. Kaur et al. [92] have proposed a wearable device capable of tracking a driver’s emotions, including anger, stress, terror, and sadness. The intelligent system, by analyzing emotion variations, determines whether the driver has entered a subconscious state and stops the vehicle’s DC motor accordingly.
@@
Adherence to medication schedules is vital but challenging for elderly individuals with memory issues. Fortunately, the integration of IoT offers a solution to this problem, and numerous research efforts have explored using IoT to track patients’ medication compliance. In [94], a medical box was created to remind individuals to take medications, featuring three trays for different times of the day. The system also measures vital health parameters (i.e., blood pressure, temperature, blood oxygen levels, etc.) and facilitates two-way communication between patients and doctors through a mobile application.
@@
Wheelchairs are vital tools in the lives of individuals who are physically unable to move independently, providing both physiological assistance and psychological support. However, for individuals with brain damage who lack the capability to operate a wheelchair, researchers have been exploring the addition of navigation and tracking systems, with IoT playing a crucial role in these wheelchair advancements. An advanced and automated smart wheelchair was reported in a study [97], which not only monitors movement but also offers features such as an umbrella, foot mat, head mat, and obstacle detection. These innovations have significantly improved interaction with the living environment and enhanced the user’s overall experience.
@@
The application of IoT in this field is versatile and has proven effective in various areas, including cancer treatment, sports injury recovery, stroke rehabilitation, and addressing physical disabilities. For instance, an innovative smart walker was introduced in a particular study [98]. Doctors and caregivers can access the collected data through a mobile application, facilitating better monitoring and support, as this walker utilizes a multi-modal sensor to monitor the patient’s walking pattern.
@@
Regular physical activity and maintaining a high level of fitness significantly influence the quality of an individual’s life. Developers have created various applications leveraging the IoT to facilitate fitness monitoring and promote healthier lifestyles. These approaches include assessing users’ activity levels and revealing metrics, including the duration of physical activity and periods of inactivity, by utilizing smartphone accelerometer data. Nowadays, wearable fitness trackers are readily available in the market and have gained popularity as convenient devices for monitoring fitness levels, for instance, smart mats to provide insights into users’ workout routines [100] and fitness assessment and training load monitoring to optimize athletes’ hydration strategies [101].
@@
The application of IoT in the healthcare industry is incredibly diverse, extending far beyond the previously mentioned areas. There are numerous domains where IoT has already been implemented and where its potential benefits are being realized, leading to a significant increase in the adoption of healthcare IoT (HIoT) technology. For cancer treatment, IoT-based methods, for example, have emerged as powerful tools. An innovative IoT-based cancer treatment approach is introduced in a recent study that encompasses various stages, including chemotherapy and radiotherapy [102]. Additionally, this system securely stores lab test results on a cloud server, allowing physicians to monitor medication dosages and enabling remote consultations through a dedicated mobile application. Furthermore, HIoT has found applications in detecting skin lesions [103], with notable advancements in lung cancer detection through state-of-the-art ML algorithms and IoT-based systems [104, 105, 106].
IoT has revolutionized the realm of surgical training and medical procedures by creating next-generation solutions. One such development involves a surgical training framework that employs virtual reality to simulate realistic training environments. This framework also enables interaction with surgeons from around the world, fostering collaborative learning and expertise sharing [107]. Monitoring haemoglobin levels in the blood has become more accessible through portable devices equipped with photoplethysmography sensors, light-emitting diodes (LEDs), and photodiodes. These devices enable the non-invasive measurement of haemoglobin levels, enhancing healthcare monitoring and diagnosis [109].
Numerous other HIoT applications are currently in use or under research, underscoring the ongoing revolutionary impact of IoT in the field of healthcare. It is expected that the IoT will continue to drive advancements and improvements in healthcare delivery and patient outcomes.
@@ 1.6.3 Smart Agriculture and Environment
Agriculture holds a crucial position in a country’s economic progress. Various factors, including soil moisture and environmental variables like carbon dioxide levels, temperature, and humidity, can significantly impact crop yields. To enhance agricultural outcomes, it becomes essential to implement robust surveillance systems in the fields. The integration of the IoT enables efficient achievement of this goal. The following section has explored several applications in this smart agricultural domain as well as a summary of the related research work in Table 1.7.
@@
Applications Related Works Applications Related Works
Water-saving irrigation [111, 112] Diseases Monitoring [113, 114]
Animal and plant life
information monitoring
[115, 116]
Intelligent agricultural
machinery
[117, 118]
Agricultural product quality
safety and traceability
[119, 120]
Crop growth environment
monitoring
[121, 122]
@@
Water scarcity in agriculture is a growing concern, necessitating a dynamic irrigation approach due to varying crop water requirements. IoT integration revolutionizes traditional flood irrigation, offering a solution to water shortage problems in crop growth. Yang et al. [111] have proposed an wireless sensor network based system leveraging neural networks for water-efficient irrigation. This method enhances irrigation efficiency by minimizing human intervention and reducing wastage due to excessive drainage.
@@
Various environmental factors, including temperature, humidity, air pressure, carbon dioxide levels, soil temperature, and soil pH, play a crucial role in crop growth. IoT devices integrated into agricultural systems can sense and analyze these environmental factors, enabling remote field monitoring and the creation of an optimal farming environment tailored to these variables. Lin et al. [114] have designed a wireless environmental monitoring system that harnesses soil energy to enable cost-effective remote monitoring of farmland environments.
@@
Effective agricultural production requires comprehensive monitoring of both plant and animal information, which is crucial for enhancing production, increasing profitability, and ensuring high-quality product development.
@@
    Animal Life Information Monitoring: Monitoring diverse aspects of animal behavior, including their food consumption, body temperature, activity levels, and health status, enables the tracking of their physiological and nutritional well-being, ensuring their healthy development. In [115], the authors have proposed an infrared-based body temperature measurement system for pigs, allowing the early detection of diseases.
@@
    Plant Life Information Monitoring: Wireless sensor devices, when connected to plants, enable remote and continuous monitoring of both external factors (such as diseases, pests, and leaf color) and internal factors (including chlorophyll content and photosynthetic rate). This technology allows for early disease detection and promotes overall healthy plant growth. Porto et al. [116] have introduced a citrus traceability system that assesses environmental conditions for optimal growth, identifying and preventing plant diseases to ensure robust citrus crop health.
@@
Intelligent machinery autonomously manages a wide range of agricultural operations, such as cultivation, sowing, transplanting, fertilization, pesticide application, feeding, irrigation, picking, and harvesting, all executed with precision and efficiency. Moreover, it has the capability to gather a variety of data about the farm, including soil moisture and water quality, as well as ambient information like temperature and humidity, which can be effectively harnessed to implement precision agriculture and enhance breeding practices [123]. IoT technology plays a key role in minimizing manual labor in agriculture by enabling remote monitoring and standardizing machinery functions through sensors and wireless communication [123]. Sowjanya et al. have introduced a versatile autonomous robot vehicle in [117] where the vehicle is equipped with Bluetooth technology for remote control and is capable of independently executing a range of tasks including farming, seeding, and irrigation.
@@
IoT significantly improves agricultural product quality, safety, and traceability, particularly in warehousing, logistics, and distribution, by enabling automatic identification, tracking, and accurate tallying of products. Various countries have implemented real-time traceability systems, such as the American, European, Swedish, Japanese, and Australian systems, recognizing the crucial need for effective tracking in agriculture. Jiang et al. [119] developed a comprehensive agricultural product safety traceability platform, facilitating real-time automatic data collection, processing, and display to enhance traceability and reduce associated tracking costs.
@@
The integration of IoT for real-time and continuous monitoring offers farmers and relevant authorities the capability to identify diseases at an early stage and implement preventive measures before they escalate. The farm’s environment plays a pivotal role in disease occurrence. For example, the framework introduced in [121] integrates various sensor devices through wireless sensor networks for monitoring various environmental factors.